# Return of the Lepton Number: Sterile Neutrino Dark Matter Production and the Revival of the Shi-Fuller Mechanism

### v1

@@ Abstract
We explore resonant production of sterile neutrino dark matter via the Shi-Fuller (SF) mechanism, revisiting its cosmological viability in light of recent results demonstrating that lepton-number asymmetries $L_{\alpha}\gtrsim 1$ at temperatures $T>20\rm\,MeV$ are consistent with Big Bang Nucleosynthesis (BBN). Using a quasi-classical Boltzmann transport calculation of the dark matter production, we compute the non-thermal phase space distributions of sterile neutrinos across a broad range of particle mass $m_{s}$ and mixing angle $\sin^{2}{(2\theta)}$ parameter space. We then evolve the resulting distributions through linear structure formation using CLASS and fit the resulting matter power spectra to thermal warm dark matter (WDM) transfer functions, enabling a direct mapping between SF models and equivalent thermal WDM particle masses $m_{\mathrm{th}}$. This allows us to reinterpret existing structure formation limits and Lyman-$\alpha$ forest preferences in the context of SF production. We find that lepton asymmetries $L\gtrsim 0.5$ at high temperatures open significant viable parameter space in the $m_{s}\gtrsim 10\,\mathrm{keV}$ and $\sin^{2}(2\theta)\lesssim 10^{-14}$ regime, compatible with both X-ray constraints from NuSTAR and INTEGRAL/SPI and recent Lyman-$\alpha$ inferences of $m_{\mathrm{th}}\approx 4.1\,\mathrm{keV}$. Following lepton number evolution below 20 MeV, we also specifically show that this lepton asymmetry parameter space is compatible with BBN and cosmic microwave background constraints. We present updated constraints, a refined $m_{\mathrm{th}}$ fitting function, and power-law approximations for $L$ across the parameter space. Our results motivate future X-ray observations targeting the $\sim 20$ keV photon regime and testing of the $m_{\mathrm{th}}\gtrsim 10\,\mathrm{keV}$ WDM region.
@@ I Introduction
Significant progress has been made in constraining a wide variety of potential candidates for the cosmological dark matter, yet the particle composition of it continues to remain a mystery. One promising class of candidates are sterile neutrinos, particles that are typically responsible for the generation of neutrino masses in theories beyond the Standard Model of particle physics. Sterile neutrino dark matter has also risen in interest because it can affect cosmological structure formation through its properties as warm to cold dark matter. In addition, it can be detected via X-ray astronomy and probed through laboratory nuclear decay experiments (for reviews, see Refs. [1, 2, 3, 4]).
Important clues to the dark matter problem are present in the distribution and properties of galaxies in the nearby Universe, and the clustering of gas and dark matter in the high-redshift Universe. Importantly, while $\Lambda$CDM—a model with a cosmological constant ($\Lambda$), and cold dark matter (CDM)—has been relatively successful despite its simple parameterization, there has been ongoing interest in extended theories of dark matter that modify its small-scale structure that could accommodate apparent discrepancies with $\Lambda$CDM (for a review, see Ref. [5]). One such simple modification is the introduction of a non-trivial initial velocity distribution for the particle dark matter from its initial thermal contact with the primordial plasma [6]. Such dark matter is now referred to as thermal warm dark matter (WDM), and its velocity dispersion results in a suppression of structure at small, sub-Galactic scales, since WDM can escape small gravitational potential wells.
@@
WDM is one of the simplest and most studied extensions of the benchmark purely cold dark matter framework. Historically, WDM was of interest in its ability to reduce the number of dwarf galaxies in the Local Group [7] as well as their central densities [8]. Now, there are a multitude of observations that constrain WDM: the phase-space density of galaxies [9, 10], satellite counts [11, 12, 13], structure in the Lyman-$\alpha$ forest [14, 15], stellar streams [16], and strong lenses [17]. The current strongest limits on thermal WDM particle masses, $m_{\mathrm{th}}$, are from galaxy counts and strong lensing, which infer $m_{\mathrm{th}}>8.4\,\mathrm{keV}$ (95% CL) [18] for galaxy counts, $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts and strong lensing [19]. Interestingly, an analysis of Lyman-$\alpha$ forest data prefers WDM over CDM, with an inferred thermal WDM particle mass of $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [20]. [^1] We have modified these $m_{\mathrm{th}}$ values to the more accurate thermal WDM transfer functions of Ref. [21], and provide this correction in Eq. (8). Within their respective analyses, using prior thermal WDM transfer functions, these works infer a thermal WDM preference of $m_{\mathrm{th}}=4.5^{+40.5}_{-1.4}\,\mathrm{keV}$ in the Lyman-$\alpha$ forest data [20] and constraints of $m_{\mathrm{th}}>9.7\,\mathrm{keV}$ (95% CL) [18] for galaxy counts, $m_{\mathrm{th}}>9.8\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts and strong lensing [19]. Sterile neutrino dark matter, however, is non-thermal, and these limits must be converted to where its particle mass produces commensurate effects [19].
@@
Another key observational property of sterile neutrino dark matter is its radiative decay channel, in which a sterile neutrino dominant mass eigenstate decays into a mass eigenstate that is predominantly an active neutrino and an X-ray photon with energies at half the dark matter particle mass [22, 23]. Although this decay rate is suppressed by the small active-sterile mixing angle, $\sin^{2}{2\theta}$, it increases as the fifth power of the particle mass, such that the signal from dark matter halos remains potentially observable with X-ray telescopes [24, 25]. Interest in sterile neutrino dark matter rose when Bulbul et al. [26] reported a high-significance, $5\sigma$, detection of an X-ray line in stacked observations of 73 clusters with the MOS and PN spectrometers aboard the XMM-Newton telescope, as well as a consistent signal from the Perseus cluster of galaxies observed with the Chandra telescope. Boyarsky et al. [27] found a consistent signal from the Andromeda galaxy as well as Perseus using data from XMM-Newton. There has been significant scrutiny of the results as well as follow-up observations that see commensurate signals in other astronomical observations [4], and several observations do not see the line with deep exposures [28, 29, 30]. Among the most significant constraints are those by Dessert et al. [31]; however, they are a factor of $\sim$20 weaker than claimed, which is acknowledged within that work, and in subsequent comments, e.g., Refs. [32, 33]. Other strong limits from Foster et al. [34] do not include instrumental and on-sky lines present at 3.3 and 3.7 keV in their stated constraints. Dessert et al. [35] claims that background mismodeling led to the original line detection, though this remains disputed [36]. The recently launched XRISM Observatory will have great sensitivity to this line [37]. The XRISM Observatory’s factor of $\sim$50 greater energy resolution than Chandra or XMM-Newton could also potentially resolve the dark matter velocity broadening of the line [26, 38].
@@
In this work, we evaluate sterile neutrino dark matter in light of X-ray and WDM constraints. The remaining parameter space is sensitive to observations from the NuSTAR observatory [39, 40, 41, 42] and SPI/INTEGRAL [43]. NuSTAR has placed leading limits on sterile neutrino decay through multiple analyses: deep “0-bounce” observations of the Galactic Center [39], pointed observations of the Bullet Cluster [40], a recent long-exposure, stacked-field survey of the Milky Way halo totaling over 20 Ms [41], and a full-mission stray-light analysis with robust background filtering spanning 11 years of data [42]. These studies find no evidence of monochromatic decay lines and set 95% C.L. upper limits at $\sin^{2}(2\theta)\lesssim 10^{-12}$ in the 6–40 keV mass range. SPI, the high-resolution spectrometer aboard INTEGRAL, searched for decay signatures in stacked galaxy cluster fields and across the Galactic halo, similarly reporting null results and exclude lifetimes below $\sim 10^{20}$ s at 95% C.L. (corresponding to $\sin^{2}(2\theta)\lesssim 3\times 10^{-14}$) in the $\gtrsim 40$ keV mass regime [43].
Many particle dark matter models propose to utilize thermal production mechanisms wherein an initial period of coupling to the standard model plasma is followed by a freeze-out which locks in the dark matter density and preserves a Fermi-Dirac phase-space distribution. Such freeze-out is also the case for early classes of WDM models [6, 44]. In contrast, WDM can be produced by a different set of mechanisms resulting in phase-space distributions that diverge from thermal for the resulting dark matter. For example, sterile neutrinos are generally produced with a non-thermal distribution, resulting from neutrino oscillations and scattering in the early Universe. In the original mechanism introduced by Dodelson & Widrow [45] (hereafter DW), active neutrinos are suppressed in their matter-affected mixing with sterile neutrinos at high-temperature, followed by production scattering rates falling with temperature, through what is often described as a freeze-in process that achieves the proper dark matter abundance. In the minimal case, the only new parameters in the DW mechanism are a (predominantly) sterile neutrino particle mass and a mixing angle with its (predominantly) active neutrino partner [45]. If there exists a lepton-number asymmetry in one or more flavors of active neutrinos in the early Universe [46], during dark matter production, the neutrinos’ matter-affected potential is modified, introducing the possibility of a Mikheev-Smirnov-Wolfenstein (MSW) resonance during production via the Shi-Fuller (SF) mechanism [47]. The SF mechanism can also be embedded in more extended models, e.g. $\nu$MSM [48, 49]. Due to the resonance, SF allows for smaller mixing angles than DW in producing the observed relic dark matter density. While the original DW mechanism is ruled-out as producing the entirety of the dark matter due to its narrow band in mass and mixing angle parameter space [50], adding a lepton-number asymmetry (SF) can provide more flexibility with respect to X-ray and structure formation constraints.
@@
The SF mechanism was thought to be highly constrained at small mixing angles since larger lepton numbers were required to produce the proper dark matter density for those mixing scales [51, 52, 12]. The lepton number or asymmetry in a neutrino flavor $\alpha$ is defined to be
@@
with $n_{\nu_{\alpha}}$ ($\bar{n}_{\nu_{\alpha}}$) the $\alpha$ flavor (anti)neutrino number density, and $n_{\gamma}$ the photon number density. The limit on the lepton number from primordial nucleosynthesis and assumed equilibration of neutrino flavors from oscillations was determined to be of order $L_{\alpha}\lesssim 0.05$ [53]. However, it was recently calculated that neutrino flavor oscillation physics locks in lepton numbers to potentially large and unequal values at early times, so that lepton numbers of order $L_{\alpha}\sim 10$ or larger can exist during SF dark matter production, a more than two orders of magnitude relaxation of constraints on the SF mechanism [54, 55]. This allows for much smaller mixing angles to produce the requisite dark matter density, with some caveats as to the particle production with such large $L_{\alpha}$, which we discuss below. Some initial work in this extended parameter space was performed in Ref. [56].
We also study the impacts of SF production on the inferred clustering of the dark matter with respect to WDM limits. Although the production mechanism of the sterile neutrino WDM candidates considered here is non-thermal, a key observation we make and go on to utilize is that the resulting linear matter power spectra can be well-approximated with those of thermally produced WDM models. We show that both thermal WDM and SF sterile neutrino dark matter have a similar shape to their characteristic suppression of small-scale power due to free-streaming effects. This similarity allows us to use the familiar transfer function typical of WDM to map a thermal particle mass onto each point in the SF parameter space.
In exploring the SF mechanism to produce the dark matter consistent with all constraints, we find that the parameter space for the SF mechanism is significantly broader than that inferred in previous work, primarily due to the relaxation of the possible lepton numbers present during SF sterile neutrino dark matter production. We examine the physics of resonant production in this mechanism in this broader parameter space. Given limits from X-ray astronomy and structure formation, we present where the SF mechanism remains viable. We also present the SF parameter space commensurate with the inferred preference for WDM from Villasenor et al. [20].
@@ II Methods
Figure 1: Shown in the left panel is the relevant average $q_{\mathrm{avg}}\equiv\langle p/T\rangle$ for the region of parameter space examined in the remainder of the work. In the right panels are the PSDs’ evolution as a function of temperature associated with two points on a constant sterile mass contour (corresponding to A & B in the left panel), demonstrating the sweeping of the resonance creating a hotter population at smaller mixing angles that require larger lepton asymmetries. The distributions lower their normalizations as the temperature decreases due to the production code simultaneously tracking cooling and dilution of the dark matter during production. Most dilution and cooling occurs through the QCD transition at $T=170\rm\,MeV$, which is most readily seen in case A.
@@ II.1 Dark Matter Production
In the SF mechanism, production of sterile neutrinos occurs through scattering during active-sterile neutrino oscillations in the presence of a primordial lepton asymmetry $L_{\alpha}$ that enhances production via a MSW-type resonance in the early Universe [47]. This production can match the observed dark matter density for smaller mixing angles than for the DW mechanism, and produces nonthermal PSDs that are colder (or “cooler”) than thermal, i.e. $\langle q\rangle\lesssim 3.15$, where $q\equiv p/T$. In this work, we utilize the sterile-dm code created by Venumadhav et al. [57]. The sterile-dm code implements a detailed calculation of resonant sterile neutrino production in the early Universe via the SF mechanism, via a single sterile neutrino species, of variable mass and mixing angle, that mixes with the muon channel.
@@
We will refer to the lepton number asymmetry in the muon flavor simply as the lepton number $L$, i.e. $L\equiv L_{\mu}$. The production is computed using a collision-dominated Boltzmann formalism appropriate for temperatures $T\gtrsim 100\,\mathrm{MeV}$, where the sterile neutrino distribution function evolves under quantum-damped, non-equilibrium dynamics, with source terms depending on the effective mixing angle in the medium, the thermal potential $V_{T}$, and the asymmetry potential $V_{L}$ [57]. Between $T\approx 100\,\mathrm{MeV}$ and $T\approx 20\,\mathrm{MeV}$, very little to no dark matter production occurs in the parameter space of interest here, so that the collision-dominated Boltzmann formalism for production is sufficient.
@@
This implementation accounts for key plasma effects that influence sterile neutrino production. Namely, it includes a detailed calculation of the redistribution of the primordial lepton asymmetry across charged leptons, neutrinos, and hadrons through weak interactions, matched across the QCD transition using lattice QCD susceptibilities. The neutrino opacities are computed from first principles, incorporating tree-level leptonic and hadronic scattering processes, including meson decays below the QCD transition and quark-level processes above it. These parameters are used to solve the Boltzmann equation for the sterile neutrino PSDs. These PSDs are prerequisite for calculating cosmological structure formation in the linear regime, via cosmological perturbation and Boltzmann calculations. The production code outputs the PSDs across the production period and the state data of the run, providing $\Omega_{\mathrm{DM}}h^{2}$ and $L$ at each temperature, down to 10 MeV. While the final $\Omega_{\mathrm{DM}}h^{2}$ is provided as an input and taken to be the full dark matter density, the corresponding $L$ is found by sterile-dm via a root-finding process. The code calculates scenarios for positive $L$, as production is symmetric with respect to positive or negative lepton number, with resonance occurring for anti-neutrinos in the case of negative $L$, with otherwise identical evolution. [^2] The full code is available at: https://github.com/ntveem/sterile-dm.
@@
A key feature of SF production is that the resonance condition evolves with temperature, sweeping across neutrino momentum space from low to high momenta as the Universe expands and cools, preferring production of low momenta modes in some parts of the mechanism’s parameter space. Much of the physics of SF production is explored in Ref. [52]. From Ref. [24], the position of the resonance is
@@
These combined effects during production are shown in Figure 1, where two PSDs with the same sterile neutrino mass exhibit different average momenta. The case with lower mixing—and thus higher lepton asymmetry—is visibly warmer. Therefore, the PSDs go from warm near the DW mechanism (small $L$), to colder at intermediate $L$, to warmer at the highest $L$ values. For a fixed sterile neutrino dark matter particle mass, this warm–cold–warm evolution in the PSD is the key result from how lepton asymmetry affects production in the SF mechanism.
@@
There are limitations to the calculation of production that we employ, which can be placed in two categories: first, there are approximations which break down as $L$ approaches and surpasses unity; and, second, the approximations of the two-neutrino mixing treatment of production via a quasi-classical Boltzmann evolution. Regarding the first category, while exploring where the SF mechanism can produce the requisite dark matter density, we therefore push the code to larger sterile neutrino masses and smaller mixing angles, extending to large $L\sim 10$ lepton asymmetries. In the production calculations, Fermi-blocking due to large $L$ is present in the quasi-classical Boltzmann evolution equations. However, the presence of large $L$ is not included in the calculations of the scattering cross-sections. Therefore, our work at $L\gtrsim 1$ is approximate and meant to be a first exploration of this parameter space.
@@
The second category of approximations include the assumption of one flavor carrying all the asymmetry, collisional decoherence treated by a simple damping term, and linear perturbative redistribution of the asymmetry in the plasma. Moreover, since lepton numbers can be redistributed due to scattering effects as well as neutrino oscillations, a more accurate calculation would solve the full quantum kinetic equations in four flavors: the sterile neutrino as well as the three active flavors, which can all carry lepton number and could all mix with each other. Such conditions can produce multiple MSW level crossings and, in the case of multiple sterile neutrinos, conditions could exist for lepton number generation [58, 59, 60]. The effects of large $L$ also complicate this second category of approximations. We leave these more thorough calculations for future work. Given these caveats at $L\gtrsim 1$, our analysis below provides the contours where $L$ becomes large in order to highlight the region where a more detailed approach is potentially necessary.

### v3

@@ Abstract
We explore resonant production of sterile neutrino dark matter via the Shi-Fuller (SF) mechanism, revisiting its cosmological viability in light of recent results demonstrating that lepton-number asymmetries $L_{\alpha}\gtrsim 1$ at temperatures $T>20\rm\,MeV$ are consistent with big bang nucleosynthesis (BBN). Using a quasiclassical Boltzmann transport calculation of the dark matter production, we compute the nonthermal phase space distributions of sterile neutrinos across a broad range of particle mass $m_{s}$ and mixing angle $\sin^{2}{(2\theta)}$ parameter space. We then evolve the resulting distributions through linear structure formation using CLASS and fit the resulting matter power spectra to thermal warm dark matter (WDM) transfer functions, enabling a direct mapping between SF models and equivalent thermal WDM particle masses $m_{\mathrm{th}}$. This allows us to reinterpret existing structure formation limits and Lyman-$\alpha$ forest preferences in the context of SF production. We find that lepton asymmetries $L\gtrsim 0.5$ at high temperatures open significant viable parameter space in the $m_{s}\gtrsim 10\,\mathrm{keV}$ and $\sin^{2}(2\theta)\lesssim 10^{-14}$ regime, compatible with both x-ray constraints from NuSTAR and INTEGRAL/SPI and recent Lyman-$\alpha$ inferences of $m_{\mathrm{th}}\approx 4.1\,\mathrm{keV}$. Following lepton number evolution below 20 MeV, we also specifically show that this lepton asymmetry parameter space is compatible with BBN and cosmic microwave background constraints. We present updated constraints, a refined $m_{\mathrm{th}}$ fitting function, and power-law approximations for $L$ across the parameter space. Our results motivate future x-ray observations targeting the $\sim 20$ keV photon regime and testing of the $m_{\mathrm{th}}\gtrsim 10\,\mathrm{keV}$ WDM region.
@@ I Introduction
Significant progress has been made in constraining a wide variety of potential candidates for the cosmological dark matter, yet the particle composition of it continues to remain a mystery. One promising class of candidates is sterile neutrinos, particles that are typically responsible for the generation of neutrino masses in theories beyond the Standard Model of particle physics. Sterile neutrino dark matter has also risen in interest because it can affect cosmological structure formation through its properties as warm to cold dark matter. In addition, it can be detected via x-ray astronomy and probed through laboratory nuclear decay experiments (for reviews, see Refs. [Boyarsky:2018tvu, Kusenko:2009up, Drewes:2016upu, Abazajian:2017tcc]).
[…]
@@
WDM is one of the simplest and most studied extensions of the benchmark purely cold dark matter framework. Historically, WDM was of interest in its ability to reduce the number of dwarf galaxies in the Local Group [Bode:2000gq] as well as their central densities [Avila-Reese:2000nqd]. Now, there are a multitude of observations that constrain WDM: the phase-space density of galaxies [Wang:2017hof, Bezrukov:2025ttd], satellite counts [Polisensky:2010rw, Cherry:2017dwu, Nadler:2019zrb], structure in the Lyman-$\alpha$ forest [Irsic:2017ixq, Palanque-Delabrouille:2019iyz], stellar streams [Banik:2019smi], and strong lenses [Keeley:2024brx]. The current strongest limits on thermal WDM particle masses, $m_{\mathrm{th}}$, are from galaxy counts and strong lensing, which infer $m_{\mathrm{th}}>8.4\,\mathrm{keV}$ (95% CL) [Nadler:2021dft] for galaxy counts, $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts and strong lensing [Zelko:2022tgf]. Interestingly, an analysis of Lyman-$\alpha$ forest data prefers WDM over CDM, with an inferred thermal WDM particle mass of $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [Villasenor:2022aiy]. [^1] We have modified these $m_{\mathrm{th}}$ values to the more accurate thermal WDM transfer functions of Ref. [Vogel:2022odl], and provide this correction in Eq. (8). Within their respective analyses, using prior thermal WDM transfer functions, these works infer a thermal WDM preference of $m_{\mathrm{th}}=4.5^{+40.5}_{-1.4}\,\mathrm{keV}$ in the Lyman-$\alpha$ forest data [Villasenor:2022aiy] and constraints of $m_{\mathrm{th}}>9.7\,\mathrm{keV}$ (95% CL) [Nadler:2021dft] for galaxy counts, $m_{\mathrm{th}}>9.8\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts and strong lensing [Zelko:2022tgf]. Sterile neutrino dark matter, however, is nonthermal, and these limits must be converted to where its particle mass produces commensurate effects [Zelko:2022tgf].
@@
Another key observational property of sterile neutrino dark matter is its radiative decay channel, in which a sterile neutrino dominant mass eigenstate decays into a mass eigenstate that is predominantly an active neutrino and an x-ray photon with energies at half the dark matter particle mass [Shrock:1974nd, Pal:1981rm]. Although this decay rate is suppressed by the small active-sterile mixing angle, $\sin^{2}{2\theta}$, it increases as the fifth power of the particle mass, such that the signal from dark matter halos remains potentially observable with x-ray telescopes [Abazajian:2001nj, Abazajian:2001vt]. Interest in sterile neutrino dark matter rose when Bulbul et al. [Bulbul:2014sua] reported a high significance, $5\sigma$, detection of an x-ray line in stacked observations of 73 clusters with the MOS and PN spectrometers aboard the XMM-Newton telescope, as well as a consistent signal from the Perseus cluster of galaxies observed with the Chandra telescope. Boyarsky:2014jta found a consistent signal from the Andromeda galaxy as well as Perseus using data from XMM-Newton. There has been significant scrutiny of the results as well as follow-up observations that see commensurate signals in other astronomical observations [Abazajian:2017tcc], and several observations do not see the line with deep exposures [Ruchayskiy:2015onc, Jeltema:2015mee, Sicilian:2020glg]. Among the most significant constraints are those by Dessert:2018qih; however, they are a factor of $\sim$20 weaker than claimed, which is acknowledged within that work, and in subsequent comments, e.g., Refs. [Boyarsky:2020hqb, Abazajian:2020unr]. Other strong limits from Foster et al. [Foster:2021ngm] do not include instrumental and on-sky lines present at 3.3 and 3.7 keV in their stated constraints. Dessert:2023fen claim that background mismodeling led to the original line detection, though this remains disputed [BulbulPrivate]. The recently launched XRISM Observatory will have great sensitivity to this line [Dessert:2023vyl]. The XRISM Observatory’s factor of $\sim$50 greater energy resolution than Chandra or XMM-Newton could also potentially resolve the dark matter velocity broadening of the line [Bulbul:2014sua, Lovell:2023olv].
@@
In this work, we evaluate sterile neutrino dark matter in light of x-ray and WDM constraints. The remaining parameter space is sensitive to observations from the NuSTAR observatory [Perez:2016tcq, Roach:2019ctw, Roach:2022lgo, Krivonos:2024yvm] and SPI/INTEGRAL [Fischer:2022pse]. NuSTAR has placed leading limits on sterile neutrino decay through multiple analyses: deep “0-bounce” observations of the Galactic Center [Perez:2016tcq], pointed observations of the Bullet Cluster [Roach:2019ctw], a recent long-exposure, stacked-field survey of the Milky Way halo totaling over 20 Ms [Roach:2022lgo], and a full-mission stray-light analysis with robust background filtering spanning 11 years of data [Krivonos:2024yvm]. These studies find no evidence of monochromatic decay lines and set 95% CL upper limits at $\sin^{2}(2\theta)\lesssim 10^{-12}$ in the 6–40 keV mass range. SPI, the high-resolution spectrometer aboard INTEGRAL, searched for decay signatures in stacked galaxy cluster fields and across the Galactic halo, similarly reporting null results and exclude lifetimes below $\sim 10^{20}$ s at 95% CL [corresponding to $\sin^{2}(2\theta)\lesssim 3\times 10^{-14}$] in the $\gtrsim 40$ keV mass regime [Fischer:2022pse].
[…]
@@
The SF mechanism was thought to be highly constrained at small mixing angles since larger lepton numbers were required to produce the proper dark matter density for those mixing scales [Chu:2006ua, Kishimoto:2008ic, Cherry:2017dwu]. The lepton number or asymmetry in a neutrino flavor $\alpha$ is defined to be
@@
with $n_{\nu_{\alpha}}$ ($\bar{n}_{\nu_{\alpha}}$) the $\alpha$ flavor (anti)neutrino number density, and $n_{\gamma}$ the photon number density. The limit on the lepton number from primordial nucleosynthesis and assumed equilibration of neutrino flavors from oscillations was determined to be of order $L_{\alpha}\lesssim 0.05$ [Serpico:2005bc]. However, it was recently calculated that neutrino flavor oscillation physics locks in lepton numbers to potentially large and unequal values at early times, so that lepton numbers of order $L_{\alpha}\sim 10$ or larger can exist during SF dark matter production, a more than 2 orders of magnitude relaxation of constraints on the SF mechanism [Froustey:2024mgf, Domcke:2025lzg]. This allows for much smaller mixing angles to produce the requisite dark matter density, with some caveats as to the particle production with such large $L_{\alpha}$, which we discuss below. Some initial work in this extended parameter space was performed in Ref. [Gorbunov:2025nqs].
We also study the impacts of SF production on the inferred clustering of the dark matter with respect to WDM limits. Although the production mechanism of the sterile neutrino WDM candidates considered here is nonthermal, a key observation we make and go on to utilize is that the resulting linear matter power spectra can be well approximated with those of thermally produced WDM models. We show that both thermal WDM and SF sterile neutrino dark matter have a similar shape to their characteristic suppression of small-scale power due to free-streaming effects. This similarity allows us to use the familiar transfer function typical of WDM to map a thermal particle mass onto each point in the SF parameter space.
In exploring the SF mechanism to produce the dark matter consistent with all constraints, we find that the parameter space for the SF mechanism is significantly broader than that inferred in previous work, primarily due to the relaxation of the possible lepton numbers present during SF sterile neutrino dark matter production. We examine the physics of resonant production in this mechanism in this broader parameter space. Given limits from x-ray astronomy and structure formation, we present where the SF mechanism remains viable. We also present the SF parameter space commensurate with the inferred preference for WDM from Villasenor et al. [Villasenor:2022aiy].
@@ II Methods
Figure 1: Shown in the left panel is the relevant average $q_{\mathrm{avg}}\equiv\langle p/T\rangle$ for the region of parameter space examined in the remainder of the work. In the right panels are the PSDs’ evolution as a function of temperature associated with two points on a constant sterile mass contour (corresponding to A and B in the left panel), demonstrating the sweeping of the resonance creating a hotter population at smaller mixing angles that require larger lepton asymmetries. The distributions lower their normalizations as the temperature decreases due to the production code simultaneously tracking cooling and dilution of the dark matter during production. Most dilution and cooling occurs through the QCD transition at $T=170\rm\,MeV$, which is most readily seen in case A.
@@ II.1 Dark matter production
In the SF mechanism, production of sterile neutrinos occurs through scattering during active-sterile neutrino oscillations in the presence of a primordial lepton asymmetry $L_{\alpha}$ that enhances production via a MSW-type resonance in the early Universe [Shi:1998km]. This production mechanism can reproduce the observed dark matter density using smaller mixing angles than those required in the DW mechanism, and produces nonthermal PSDs that are colder (or “cooler”) than thermal, i.e., $\langle q\rangle\lesssim 3.15$, where $q\equiv p/T$. In this work, we utilize the sterile-dm code created by Venumadhav et al. [Venumadhav:2015pla]. The sterile-dm code implements a detailed calculation of resonant sterile neutrino production in the early Universe via the SF mechanism, via a single sterile neutrino species, of variable mass and mixing angle, that mixes with the muon channel.
@@
We will refer to the lepton number asymmetry in the muon flavor simply as the lepton number $L$, i.e. $L\equiv L_{\mu}$. The production is computed using a collision-dominated Boltzmann formalism appropriate for temperatures $T\gtrsim 100\,\mathrm{MeV}$, where the sterile neutrino distribution function evolves under quantum-damped, nonequilibrium dynamics, with source terms depending on the effective mixing angle in the medium, the thermal potential $V_{T}$, and the asymmetry potential $V_{L}$ [Venumadhav:2015pla]. Between $T\approx 100\,\mathrm{MeV}$ and $T\approx 20\,\mathrm{MeV}$, very little to no dark matter production occurs in the parameter space of interest here, so that the collision-dominated Boltzmann formalism for production is sufficient.
@@
This implementation accounts for key plasma effects that influence sterile neutrino production. Namely, it includes a detailed calculation of the redistribution of the primordial lepton asymmetry across charged leptons, neutrinos, and hadrons through weak interactions, matched across the QCD transition using lattice QCD susceptibilities. The neutrino opacities are computed from first principles, incorporating tree-level leptonic and hadronic scattering processes, including meson decays below the QCD transition and quark-level processes above it. These parameters are used to solve the Boltzmann equation for the sterile neutrino PSDs. These PSDs are prerequisite for calculating cosmological structure formation in the linear regime, via cosmological perturbation and Boltzmann calculations. The production code outputs the PSDs across the production period and the state data of the run, providing $\Omega_{\mathrm{DM}}h^{2}$ and $L$ at each temperature, down to 10 MeV. While the final $\Omega_{\mathrm{DM}}h^{2}$ is provided as an input and taken to be the full dark matter density, the corresponding $L$ is found by sterile-dm via a root-finding process. The code calculates scenarios for positive $L$, as production is symmetric with respect to positive or negative lepton number, with resonance occurring for antineutrinos in the case of negative $L$, with otherwise identical evolution. [^2] The full code is available at: https://github.com/ntveem/sterile-dm.
@@
A key feature of SF production is that the resonance condition evolves with temperature, sweeping across neutrino momentum space from low to high momenta as the Universe expands and cools, preferring production of low momenta modes in some parts of the mechanism’s parameter space. Much of the physics of SF production is explored in Ref. [Kishimoto:2008ic]. From Ref. [Abazajian:2001nj], the position of the resonance is
@@
These combined effects during production are shown in Fig.1, where two PSDs with the same sterile neutrino mass exhibit different average momenta. The case with lower mixing—and thus higher lepton asymmetry—is visibly warmer. Therefore, the PSDs go from warm near the DW mechanism (small $L$), to colder at intermediate $L$, to warmer at the highest $L$ values. For a fixed sterile neutrino dark matter particle mass, this warm–cold–warm evolution in the PSD is the key result from how lepton asymmetry affects production in the SF mechanism.
@@
There are limitations to the calculation of production that we employ, which can be placed in two categories: first, there are approximations which break down as $L$ approaches and surpasses unity; and, second, the approximations of the two-neutrino mixing treatment of production via a quasiclassical Boltzmann evolution. Regarding the first category, while exploring where the SF mechanism can produce the requisite dark matter density, we therefore push the code to larger sterile neutrino masses and smaller mixing angles, extending to large $L\sim 10$ lepton asymmetries. In the production calculations, Fermi blocking due to large $L$ is present in the quasiclassical Boltzmann evolution equations. However, the presence of large $L$ is not included in the calculations of the scattering cross sections. Therefore, our work at $L\gtrsim 1$ is approximate and meant to be a first exploration of this parameter space.
@@
The second category of approximations includes the assumption of one flavor carrying all the asymmetry, collisional decoherence treated by a simple damping term, and linear perturbative redistribution of the asymmetry in the plasma. Moreover, since lepton numbers can be redistributed due to scattering effects as well as neutrino oscillations, a more accurate calculation would solve the full quantum kinetic equations in four flavors: the sterile neutrino as well as the three active flavors, which can all carry lepton number and could all mix with each other. Such conditions can produce multiple MSW level crossings and, in the case of multiple sterile neutrinos, conditions could exist for lepton number generation [Foot:1995bm, Shi:1996ic, Abazajian:2004aj]. The effects of large $L$ also complicate this second category of approximations. We leave these more thorough calculations for future work. Given these caveats at $L\gtrsim 1$, our analysis below provides the contours where $L$ becomes large in order to highlight the region where a more detailed approach is potentially necessary.
In addition, we note that previous work has found that increasing the momentum bin count in sterile-dm is increasingly important as the analysis moves to a higher lepton number [Bodeker:2020hbo]. It appears that part of the resonance is missed if the momentum resolution is low, resulting in less dark matter production. Practically, this manifests as a larger lepton number identified by the code if using an insufficient resolution. In order to mitigate this issue, we analyzed different regions in parameter space before completing our analysis, finding in agreement with the authors of Ref. [Bodeker:2020hbo] that 30,000 momentum bins is generally sufficient; however, for regions near nonresonant production, we use fewer bins (down to 5,000) and for very large lepton number regions we use more (up to 50,000) to balance computational efficiency with the maintenance of accuracy at the single-percent level ($\approx 1\%$). Figure 2 shows the results of our analysis for the two points A and B from Fig. 1. The results demonstrate the necessity for higher resolution at relatively large lepton number (smaller mixing, such as point B) but the lesser importance in the low lepton number (larger mixing, such as point A) regime.
Figure 2: The blue (orange) curve shows the percentage error in the dark matter relic density, $\omega\equiv\Omega_{\mathrm{DM}}h^{2}$, as a function of momentum bin number in the sterile-dm code for the sterile neutrino parameters associated with point A (B) in Fig. 1.